# Toward Black–Scholes for Prediction Markets: A Unified Kernel and Market-Maker’s Handbook

### v1

@@ 1 Introduction
Prediction markets like Polymarket trade contracts that pay \$1 if an event occurs and \$0 otherwise. Under standard no-arbitrage reasoning, observed prices are interpretable as risk-neutral probabilities of the event. Empirically, these prices often track average beliefs under mild conditions, although biases exist and interpretation requires care (WolfersZitzewitz2004, ; WolfersZitzewitz2006, ; Manski2004, ). These markets have been used to forecast outcomes in politics, economics, science, and beyond (Arrow2008, ; Snowberg2012, ). Beyond entertainment or wagering, liquid, global event markets instantiate a Hayekian mechanism for aggregating dispersed knowledge (Hayek1945, ). Prices synthesize private signals into public probabilities, creating incentives for informed actors to reveal information when they expect to profit. Empirically and in practice, prediction markets can improve forecast accuracy and organizational decision-making by rewarding correct beliefs and penalizing noise (CowgillZitzewitz2015, ). The appeal has long been recognized by policymakers: the U.S. Defense Advanced Research Projects Agency (DARPA) explored a Policy Analysis Market (FutureMAP) for geopolitical risks before cancelling it in 2003 amid political criticism (WaPo2003, ; Hanson2007PAM, ). At the same time, regulatory choices shape what can be listed: for example, the 1958 Onion Futures Act still bans U.S. onion futures, illustrating how legal constraints can limit the scope of information aggregation via markets (USC7_13_1, ). Finally, recent debates over polling errors (e.g., the 2016 U.S. election) underscore the value of complementary, market-based signals when traditional information channels are noisy or biased (AAPOR2016, ). Together, these strands motivate a standardized kernel and derivative layer so that “belief risk” can be quoted, hedged, and transferred at scale.
@@
An event contract is a yes/no claim with a fixed payoff: it pays \$1 if the specified outcome occurs by a given date and \$0 otherwise. The traded price $p_{t}\!\in\!(0,1)$ is naturally read as the market’s risk-neutral probability for that outcome. Examples include “Will the unemployment rate exceed 5% in Q4?” or “Will upgrade $U$ activate by block height HH?”. In regulated settings, the U.S. Commodity Futures Trading Commission (CFTC) describes such claims as derivative contracts tied to the occurrence of an event, typically with binary payoff structure; exchanges listing them must satisfy the same substantive requirements as other derivatives (FedRegisterEventContracts2024, ; CFTCEventContracts, ). Crypto-native venues list economically similar contracts under on-chain execution.
Execution is fragmented. Platforms rely on the Logarithmic Market Scoring Rule (LMSR) (Hanson2003, ), on constant-product automated market makers (CP-AMMs)—a special case of constant-function market makers (CFMMs)—popularized by Uniswap (UniswapV2, ; AngerisCFMM, ), or on traditional order books. Each mechanism sets prices, but none provides a shared stochastic kernel that explains how event probabilities should evolve over time, across information shocks, or jointly across related events. By contrast, once the Black–Scholes (BS) model appeared in options, markets standardized around implied volatility, which enabled quoting, hedging, and a deep derivative layer (BlackScholes1973, ). Prediction markets lack an analogous foundation.
@@
Market making in event contracts is hard. In microstructure, informed trading induces adverse selection: when a counterparty trades only when you are wrong, you lose on average (GlostenMilgrom1985, ; Kyle1985, ). In binary event markets, inventory risk is concentrated near resolution and cannot be hedged by the underlying until settlement. For cost-function market makers such as LMSR, providing more liquidity increases worst-case loss; without external subsidies, they are expected to run at a deficit proportional to the liquidity they offer (OthmanSandholm2010, ). CP-AMMs/CFMMs face related profitability constraints for liquidity providers (AngerisUniswap2019, ; Bitterli2023, ). In short, today’s mechanisms expose makers to toxic flow and gap risk, but do not offer standardized tools to transfer belief risk (the risk that the market-implied probability moves) across time or across related events.
@@
Derivatives exist to let participants isolate and trade specific risks. For events, the primary risks are movements in the belief level and its volatility (how fast log-odds move), plus jump risk from news and cross-event co-movement. A derivative layer would (i) let market makers hedge adverse selection by offloading belief volatility and jump exposure; (ii) allow calendar hedges (between maturities or checkpoints before resolution); (iii) enable cross-event hedges that neutralize correlation and co-jumps; and (iv) concentrate liquidity around a small set of quoted risk factors, as implied volatility did for options. In mature option markets, variance/volatility swaps, correlation swaps, and related instruments serve exactly these functions for price volatility (Demeterfi1999, ; CarrLee2009, ; Broadie2008, ). Prediction markets need the analogous instruments for belief dynamics.
@@
Crypto-native platforms have brought event trading to a wider audience and catalyzed institutional interest. In 2025, Intercontinental Exchange (ICE), the parent of the New York Stock Exchange, announced a strategic investment of up to \$2 billion for roughly a 20% stake in Polymarket, with plans to distribute event-driven data through ICE’s channels (AxiosICE2025, ). Separate disclosures report that Polymarket raised over \$200 million across 2024–2025 prior to the ICE deal (TheBlock205M2025, ). Together with growth spurts around major elections and macro events, these developments signal mainstream acceptance of event markets and amplify the need for a common pricing kernel and a standardized derivative layer.
@@
Standardization matters more than perfect realism. Black–Scholes was not “true,” but it coordinated quoting and hedging around a small number of state variables; that standard enabled scale. A shared belief–variance surface can play the same role for event markets precisely as adoption accelerates. In 2025, Intercontinental Exchange (ICE), owner of the NYSE, announced up to \$2 billion of strategic investment in Polymarket (at roughly \$8 billion pre-money) and will distribute its event-driven data to institutions, a signal of mainstream integration (ICE2025PR, ; FT2025Polymarket, ; Reuters2025Polymarket, ). On the usage side, monthly volumes have broken through billion-dollar thresholds: reports based on Dune/DeFiLlama data indicate Polymarket posted about \$1.4 billion in September 2025 while Kalshi exceeded \$1–3 billion depending on the week and product mix, with sports driving much of the surge (Defiant2025PM, ; Decrypt2025Kalshi, ; LSR2025Kalshi, ). At the same time, U.S. oversight is becoming more explicit: the CFTC proposed rulemaking on event contracts in 2024 (clarifying categories that may be contrary to the public interest), and litigation around political markets underscores an evolving but increasingly articulate framework (CFTC2024NPRM, ; FR2024EventContracts, ; KalshiCADC2024, ). This combination of institutional capital, record volumes, and regulatory clarity strengthens the case for a common pricing kernel and a standardized derivative layer to concentrate liquidity, reduce maker losses via hedging, and support an institutional market for belief risk.
@@ 2 Related Work
The Logarithmic Market Scoring Rule (LMSR) introduced a bounded-loss, always-on automated market maker (AMM) for event contracts and combinatorial claims, giving crisp axiomatic guarantees and a tractable cost function representation (Hanson2003, ; ChenPennock2012, ). Subsequent designs generalized AMMs to convex cost-function markets and to liquidity-adaptive variants that mitigate worst-case losses and over-movement in thin markets (Abernethy2013, ; OthmanSandholm2010, ). On crypto rails, constant-function market makers (CFMMs)—including the constant-product AMM (CP-AMM) popularized by Uniswap—standardized execution and on-chain pricing primitives (AngerisChitra2020, ; UniswapV2, ). In contrast, central limit order books (CLOBs) provide deep execution microstructure but no common probabilistic dynamics for event prices (Gould2013, ). Our focus is complementary: we seek a shared stochastic kernel for risk-neutral event probabilities across time, shocks, and related events, independent of the execution venue.
@@
The Black–Scholes–Merton paradigm established a common language for quoting and hedging (implied volatility, Greeks) (BlackScholes1973, ). It catalyzed successive layers: jumps (Merton1976, ), stochastic volatility (Heston1993, ), and implied-consistent local volatility surfaces (Dupire1994, ), leading to the modern practice of surface construction and management (Gatheral2006, ). This standardization coordinated liquidity and risk transfer. Our work aims for the analogous role in prediction markets: replace price diffusions by logit dynamics for probabilities on $(0,1)$, preserving tractability while exposing belief-level, belief-volatility, and jump/correlation factors as quotable objects.
@@
Information-based asset pricing treats prices as conditional expectations under a filtration generated by noisy “information processes,” often Brownian-bridge–driven, offering a structural view of how news reveals payoffs over time (BrodyHughstonMacrina2007, ). Separately, boundary-constrained diffusions on $[0,1]$ (e.g., Wright–Fisher/Jacobi families) provide mathematically consistent dynamics for probabilities or bounded state variables (JenkinsSpano2017, ; Ackerer2016, ). We adopt a simple alternative: a logit map that transports $p_{t}\!\in(0,1)$ to $\mathbb{R}$ so standard semimartingale tools apply, while explicit jump terms capture news shocks and co-jumps across related events.
@@
Our calibration pipeline draws on state-space methods that separate the latent “efficient” signal from microstructure noise in high-frequency mid/bid-ask/trade streams (Hasbrouck1991, ; EngleZheng2006, ). For parameter learning with jumps, EM/likelihood filters for jump-diffusions and inference tools for detecting common jumps offer practical estimators and diagnostics (BandiNguyen2001, ; BeginBoudreault2019, ; JacodTodorov2009, ). We adapt these ingredients to log-odds increments and to co-jump screening across events, yielding a stable belief-volatility surface suitable for quoting and hedging.
@@ 3.1 Background and Motivation: From Black–Scholes to Event Probabilities
self-financing replication implies a linear pricing PDE and closed-form option values. The key output is not realism per se, but a common language for quoting and hedging: implied volatility, Greeks, and volatility surfaces (BlackScholes1973, ; Merton1973, ; Heston1993, ; Dupire1994, ; Gatheral2006, ). This language coordinates liquidity, enables standardized risk transfer, and supports a deep derivative stack.
@@
Event contracts trade binary payoffs. Their quoted prices are interpretable as discounted, risk–neutral probabilities of occurrence (WolfersZitzewitz2006, ; Arrow2008, ). Today, venues execute via scoring rules or AMMs or CLOBs, but there is no shared stochastic model for how probabilities evolve across time, shocks, or related events. Without a kernel, makers cannot isolate “belief risk” (level, volatility, jumps, co-movement) or lay it off in a standard way; spreads widen around news; inventory near the $0/1$ boundaries becomes hard to manage. A tractable kernel standardizes quoting (what to post), hedging (what to buy/sell against), and calibration (how to read data), just as implied-vol surfaces did for options.
@@
(i) State variable: BS models prices; we model probabilities. We map $p_{t}\in(0,1)$ to log-odds $x_{t}\in\mathbb{R}$ so Itô–Lévy tools apply while respecting boundaries. (ii) Martingale restriction: in BS, discounted $S_{t}$ is a martingale; here discounted $p_{t}=S(x_{t})$ is a martingale. This pins down the drift of $x_{t}$ (Eq. (3)) and leaves belief-volatility and jump features as the quotable risks. (iii) News and co-movement: event probabilities jump at information times, and related events co-move; our kernel includes both diffusive correlation and co-jumps (Eq. (4)), analogous to jumps/SV in equity models (Merton1976, ; Heston1993, ; ContTankov2004, ). (iv) Incompleteness: the binary payoff cannot be dynamically replicated by the underlying before resolution, so markets are incomplete. Derivatives on $x$ or $p$ (variance, correlation, corridor, first-passage) create targeted hedges that make inventory and adverse-selection risk manageable, echoing the role of variance and correlation swaps in equities (Demeterfi1999, ; CarrLee2009, ).
@@
Across liquid option markets, the existence of a shared surface reduced dispersion in quotes and tightened spreads (Gatheral2006, ). Prediction markets show analogous frictions: spreads and cancellations widen near scheduled news; quotes gap on unexpected announcements; correlated events move together. A belief-variance/correlation layer makes these exposures explicit and tradable, allowing makers to keep quotes live while laying off risk—precisely the coordination role BS played for options.
@@ 3.2 Kernel: Logit Jump–Diffusion with Risk–Neutral Drift
where $\sigma_{b}$ is the belief volatility. This $x$-dynamics guarantees $p_{t}=S(x_{t})\in(0,1)$ while allowing diffusive moves and news-driven jumps. The representation (1) is a standard Itô–Lévy SDE (Applebaum2009, ; OksendalSulem2005, ; ContTankov2004, ).
@@
Equations (2)–(3) ensure $p_{t}$ is a $\mathbb{Q}$-martingale; therefore, only the belief-volatility $\sigma_{b}$, the jump intensity and moments (embedded in $\nu_{t}$), and cross-event dependence (introduced below) remain as tradable risk factors (Applebaum2009, ; ContTankov2004, ).
@@
The logit map transports the bounded probability $p_{t}\!\in(0,1)$ to $\mathbb{R}$ where standard semimartingale tools apply, while the jump term allows for abrupt probability updates at news times. The martingale restriction fixes the drift of $x_{t}$ so that $S(x_{t})$ carries zero drift under $\mathbb{Q}$; informally, the “belief level” $p_{t}$ drifts only when reparameterized in $x$ to offset convexity and jump-compensation effects. This separation makes $\sigma_{b}$ and jump features economically interpretable and quotable—directly analogous to how implied variance and jump parameters are quoted in price-based models (Merton1976, ; Kou2002, ; ContTankov2004, ).

### v2

@@ 1 Introduction
Prediction markets like Polymarket trade contracts that pay \$1 if an event occurs and \$0 otherwise. Under standard no-arbitrage reasoning, observed prices are interpretable as risk-neutral probabilities of the event. Empirically, these prices often track average beliefs under mild conditions, although biases exist and interpretation requires care (Wolfers and Zitzewitz, 2004, 2006; Manski, 2004). These markets have been used to forecast outcomes in politics, economics, science, and beyond (Arrow et al., 2008; Snowberg et al., 2012). Beyond entertainment or wagering, liquid, global event markets instantiate a Hayekian mechanism for aggregating dispersed knowledge (Hayek, 1945). Prices synthesize private signals into public probabilities, creating incentives for informed actors to reveal information when they expect to profit. Empirically and in practice, prediction markets can improve forecast accuracy and organizational decision-making by rewarding correct beliefs and penalizing noise (Cowgill et al., 2015). The appeal has long been recognized by policymakers: the U.S. Defense Advanced Research Projects Agency (DARPA) explored a Policy Analysis Market (FutureMAP) for geopolitical risks before cancelling it in 2003 amid political criticism (Graham and Loeb, 2003; Hanson, 2007). At the same time, regulatory choices shape what can be listed: for example, the 1958 Onion Futures Act still bans U.S. onion futures, illustrating how legal constraints can limit the scope of information aggregation via markets (United States Code, 1958). Finally, recent debates over polling errors (e.g., the 2016 U.S. election) underscore the value of complementary, market-based signals when traditional information channels are noisy or biased (American Association for Public Opinion Research, 2017). Together, these strands motivate a standardized kernel and derivative layer so that “belief risk” can be quoted, hedged, and transferred at scale.
@@
An event contract is a yes/no claim with a fixed payoff: it pays \$1 if the specified outcome occurs by a given date and \$0 otherwise. The traded price $p_{t}\!\in\!(0,1)$ is naturally read as the market’s risk-neutral probability for that outcome. Examples include “Will the unemployment rate exceed 5% in Q4?” or “Will upgrade $U$ activate by block height HH?”. In regulated settings, the U.S. Commodity Futures Trading Commission (CFTC) describes such claims as derivative contracts tied to the occurrence of an event, typically with binary payoff structure; exchanges listing them must satisfy the same substantive requirements as other derivatives (30; 59). Crypto-native venues list economically similar contracts under on-chain execution.
Execution is fragmented. Platforms rely on the Logarithmic Market Scoring Rule (LMSR) (Hanson, 2003), on constant-product automated market makers (CP-AMMs)—a special case of constant-function market makers (CFMMs)—popularized by Uniswap (Adams et al., 2020; Angeris et al., 2023), or on traditional order books. Each mechanism sets prices, but none provides a shared stochastic kernel that explains how event probabilities should evolve over time, across information shocks, or jointly across related events. By contrast, once the Black–Scholes (BS) model appeared in options, markets standardized around implied volatility, which enabled quoting, hedging, and a deep derivative layer (Black and Scholes, 1973). Prediction markets lack an analogous foundation.
@@
Market making in event contracts is hard. In microstructure, informed trading induces adverse selection: when a counterparty trades only when you are wrong, you lose on average (Glosten and Milgrom, 1985; Kyle, 1985). In binary event markets, inventory risk is concentrated near resolution and cannot be hedged by the underlying until settlement. For cost-function market makers such as LMSR, providing more liquidity increases worst-case loss; without external subsidies, they are expected to run at a deficit proportional to the liquidity they offer (Othman et al., 2010). CP-AMMs/CFMMs face related profitability constraints for liquidity providers (Angeris et al., 2019; Bitterli and Schär, 2023). In short, today’s mechanisms expose makers to toxic flow and gap risk, but do not offer standardized tools to transfer belief risk (the risk that the market-implied probability moves) across time or across related events.
@@
Derivatives exist to let participants isolate and trade specific risks. For events, the primary risks are movements in the belief level and its volatility (how fast log-odds move), plus jump risk from news and cross-event co-movement. A derivative layer would (i) let market makers hedge adverse selection by offloading belief volatility and jump exposure; (ii) allow calendar hedges (between maturities or checkpoints before resolution); (iii) enable cross-event hedges that neutralize correlation and co-jumps; and (iv) concentrate liquidity around a small set of quoted risk factors, as implied volatility did for options. In mature option markets, variance/volatility swaps, correlation swaps, and related instruments serve exactly these functions for price volatility (Demeterfi et al., 1999; Carr and Lee, 2009; Broadie and Jain, 2008b). Prediction markets need the analogous instruments for belief dynamics.
@@
Crypto-native platforms have brought event trading to a wider audience and catalyzed institutional interest. In 2025, Intercontinental Exchange (ICE), the parent of the New York Stock Exchange, announced a strategic investment of up to \$2 billion for roughly a 20% stake in Polymarket, with plans to distribute event-driven data through ICE’s channels (Primack and McCracken, 2025). Separate disclosures report that Polymarket raised over \$200 million across 2024–2025 prior to the ICE deal (Block, 2025). Together with growth spurts around major elections and macro events, these developments signal mainstream acceptance of event markets and amplify the need for a common pricing kernel and a standardized derivative layer.
@@
Standardization matters more than perfect realism. Black–Scholes was not “true,” but it coordinated quoting and hedging around a small number of state variables; that standard enabled scale. A shared belief–variance surface can play the same role for event markets precisely as adoption accelerates. In 2025, Intercontinental Exchange (ICE), owner of the NYSE, announced up to \$2 billion of strategic investment in Polymarket (at roughly \$8 billion pre-money) and will distribute its event-driven data to institutions, a signal of mainstream integration (Intercontinental Exchange, Inc., 2025; Times, 2025; Reuters, 2025). On the usage side, monthly volumes have broken through billion-dollar thresholds: reports based on Dune/DeFiLlama data indicate Polymarket posted about \$1.4 billion in September 2025 while Kalshi exceeded \$1–3 billion depending on the week and product mix, with sports driving much of the surge (Defiant, 2025; Decrypt, 2025; Report, 2025). At the same time, U.S. oversight is becoming more explicit: the CFTC proposed rulemaking on event contracts in 2024 (clarifying categories that may be contrary to the public interest), and litigation around political markets underscores an evolving but increasingly articulate framework (U.S. Commodity Futures Trading Commission, 2024a, c; Justia, 2024). This combination of institutional capital, record volumes, and regulatory clarity strengthens the case for a common pricing kernel and a standardized derivative layer to concentrate liquidity, reduce maker losses via hedging, and support an institutional market for belief risk.
@@ 2 Related Work
The Logarithmic Market Scoring Rule (LMSR) introduced a bounded-loss, always-on automated market maker (AMM) for event contracts and combinatorial claims, giving crisp axiomatic guarantees and a tractable cost function representation (Hanson, 2003; Chen and Pennock, 2012). Subsequent designs generalized AMMs to convex cost-function markets and to liquidity-adaptive variants that mitigate worst-case losses and over-movement in thin markets (Abernethy et al., 2013; Othman et al., 2010). On crypto rails, constant-function market makers (CFMMs)—including the constant-product AMM (CP-AMM) popularized by Uniswap—standardized execution and on-chain pricing primitives (Angeris and Chitra, 2020; Adams et al., 2020). In contrast, central limit order books (CLOBs) provide deep execution microstructure but no common probabilistic dynamics for event prices (Gould et al., 2013). Our focus is complementary: we seek a shared stochastic kernel for risk-neutral event probabilities across time, shocks, and related events, independent of the execution venue.
@@
The Black–Scholes–Merton paradigm established a common language for quoting and hedging (implied volatility, Greeks) (Black and Scholes, 1973). It catalyzed successive layers: jumps (Merton, 1976), stochastic volatility (Heston, 1993), and implied-consistent local volatility surfaces (Dupire, 1994), leading to the modern practice of surface construction and management (Gatheral, 2006). This standardization coordinated liquidity and risk transfer. Our work aims for the analogous role in prediction markets: replace price diffusions by logit dynamics for probabilities on $(0,1)$, preserving tractability while exposing belief-level, belief-volatility, and jump/correlation factors as quotable objects.
@@
Information-based asset pricing treats prices as conditional expectations under a filtration generated by noisy “information processes,” often Brownian-bridge–driven, offering a structural view of how news reveals payoffs over time (Brody et al., 2007). Separately, boundary-constrained diffusions on $[0,1]$ (e.g., Wright–Fisher/Jacobi families) provide mathematically consistent dynamics for probabilities or bounded state variables (Jenkins and Spanò, 2017; Ackerer et al., 2016). We adopt a simple alternative: a logit map that transports $p_{t}\!\in(0,1)$ to $\mathbb{R}$ so standard semimartingale tools apply, while explicit jump terms capture news shocks and co-jumps across related events.
@@
Our calibration pipeline draws on state-space methods that separate the latent “efficient” signal from microstructure noise in high-frequency mid/bid-ask/trade streams (Hasbrouck, 1991; Engle and Sun, 2006). For parameter learning with jumps, EM/likelihood filters for jump-diffusions and inference tools for detecting common jumps offer practical estimators and diagnostics (Bandi and Nguyen, 2001; Bégin and Boudreault, 2019; Jacod and Todorov, 2009). We adapt these ingredients to log-odds increments and to co-jump screening across events, yielding a stable belief-volatility surface suitable for quoting and hedging.
@@ 3.1 Background and Motivation: From Black–Scholes to Event Probabilities
self-financing replication implies a linear pricing PDE and closed-form option values. The key output is not realism per se, but a common language for quoting and hedging: implied volatility, Greeks, and volatility surfaces (Black and Scholes, 1973; Merton, 1973; Heston, 1993; Dupire, 1994; Gatheral, 2006). This language coordinates liquidity, enables standardized risk transfer, and supports a deep derivative stack.
@@
Event contracts trade binary payoffs. Their quoted prices are interpretable as discounted, risk–neutral probabilities of occurrence (Wolfers and Zitzewitz, 2006; Arrow et al., 2008). Today, venues execute via scoring rules or AMMs or CLOBs, but there is no shared stochastic model for how probabilities evolve across time, shocks, or related events. Without a kernel, makers cannot isolate “belief risk” (level, volatility, jumps, co-movement) or lay it off in a standard way; spreads widen around news; inventory near the $0/1$ boundaries becomes hard to manage. A tractable kernel standardizes quoting (what to post), hedging (what to buy/sell against), and calibration (how to read data), just as implied-vol surfaces did for options.
@@
(i) State variable: BS models prices; we model probabilities. We map $p_{t}\in(0,1)$ to log-odds $x_{t}\in\mathbb{R}$ so Itô–Lévy tools apply while respecting boundaries. (ii) Martingale restriction: in BS, discounted $S_{t}$ is a martingale; here discounted $p_{t}=S(x_{t})$ is a martingale. This pins down the drift of $x_{t}$ (Eq. (3)) and leaves belief-volatility and jump features as the quotable risks. (iii) News and co-movement: event probabilities jump at information times, and related events co-move; our kernel includes both diffusive correlation and co-jumps (Eq. (4)), analogous to jumps/SV in equity models (Merton, 1976; Heston, 1993; Cont and Tankov, 2004). (iv) Incompleteness: the binary payoff cannot be dynamically replicated by the underlying before resolution, so markets are incomplete. Derivatives on $x$ or $p$ (variance, correlation, corridor, first-passage) create targeted hedges that make inventory and adverse-selection risk manageable, echoing the role of variance and correlation swaps in equities (Demeterfi et al., 1999; Carr and Lee, 2009).
@@
Across liquid option markets, the existence of a shared surface reduced dispersion in quotes and tightened spreads (Gatheral, 2006). Prediction markets show analogous frictions: spreads and cancellations widen near scheduled news; quotes gap on unexpected announcements; correlated events move together. A belief-variance/correlation layer makes these exposures explicit and tradable, allowing makers to keep quotes live while laying off risk—precisely the coordination role BS played for options.
@@ 3.2 Kernel: Logit Jump–Diffusion with Risk–Neutral Drift
where $\sigma_{b}$ is the belief volatility. This $x$-dynamics guarantees $p_{t}=S(x_{t})\in(0,1)$ while allowing diffusive moves and news-driven jumps. The representation (1) is a standard Itô–Lévy SDE (Applebaum, 2009; Øksendal and Sulem, 2005; Cont and Tankov, 2004).
@@
Equations (2)–(3) ensure $p_{t}$ is a $\mathbb{Q}$-martingale; therefore, only the belief-volatility $\sigma_{b}$, the jump intensity and moments (embedded in $\nu_{t}$), and cross-event dependence (introduced below) remain as tradable risk factors (Applebaum, 2009; Cont and Tankov, 2004).
@@
The logit map transports the bounded probability $p_{t}\!\in(0,1)$ to $\mathbb{R}$ where standard semimartingale tools apply, while the jump term allows for abrupt probability updates at news times. The martingale restriction fixes the drift of $x_{t}$ so that $S(x_{t})$ carries zero drift under $\mathbb{Q}$; informally, the “belief level” $p_{t}$ drifts only when reparameterized in $x$ to offset convexity and jump-compensation effects. This separation makes $\sigma_{b}$ and jump features economically interpretable and quotable—directly analogous to how implied variance and jump parameters are quoted in price-based models (Merton, 1976; Kou, 2002; Cont and Tankov, 2004).